# Rare-to-Frequent: Unlocking Compositional Generation Power of Diffusion Models on Rare Concepts with LLM Guidance

### v1

@@ Abstract
State-of-the-art text-to-image (T2I) diffusion models often struggle to generate rare compositions of concepts, e.g., objects with unusual attributes. In this paper, we show that the compositional generation power of diffusion models on such rare concepts can be significantly enhanced by the Large Language Model (LLM) guidance. We start with empirical and theoretical analysis, demonstrating that exposing frequent concepts relevant to the target rare concepts during the diffusion sampling process yields more accurate concept composition. Based on this, we propose a training-free approach, R2F, that plans and executes the overall rare-to-frequent concept guidance throughout the diffusion inference by leveraging the abundant semantic knowledge in LLMs. Our framework is flexible across any pre-trained diffusion models and LLMs, and can be seamlessly integrated with the region-guided diffusion approaches. Extensive experiments on three datasets, including our newly proposed benchmark, RareBench, containing various prompts with rare compositions of concepts, R2F significantly surpasses existing models including SD3.0 and FLUX by up to $28.1\%p$ in T2I alignment. Code is available at https://github.com/krafton-ai/Rare2Frequent.
@@ 1 Introduction
To thoroughly validate the efficacy of R2F, we present a new benchmark, dubbed RareBench, consisting of diverse and complex rare compositions of concepts. On RareBench and two existing compositionality benchmarks (Rassin et al., 2024, Huang et al., 2023), R2F outperforms state-of-the-art diffusion baselines, including SD3.0 and FLUX, by up to $28.1\%p$ in terms of T2I alignment accuracy. Moreover, we show that R2F can generate images of rare concepts that existing models almost always fail to generate even with careful prompt paraphrasing, showcasing the superiority of our framework in unlocking the compositional generation power of diffusion models.
@@ 3.2 Proposed Framework: R2F
Inspired by the theoretical motivation, we propose a training-free framework, called R2F, which leverages LLM to find frequent concepts relevant to rare concepts and use them in the diffusion sampling process. R2F involves a two-stage process: (i) Rare-to-frequent concept mapping that uses LLM to identify and map rare concepts and their relevant yet frequent concepts; and (ii) Alternating concept guidance that iteratively uses prompts involving either rare or frequent objects during the sampling process. Different from Theorem 3.1, we use alternating guidance because it yields more realistic images than the linear composition as it continuously denoises from the actual prompts, which is analyzed in Section 4.4. Figure 4 illustrates the overview of our proposed framework.
@@ 4.1 Experiment Setting
Implementation Details. Our R2F is flexible to an arbitrary combination of diffusion models and LLMs. By default, we use SD3.0 and GPT-4o (Achiam et al., 2023). For every inference, we set the sampling steps $T$ to $50$ and the random seed to $42$. The hyperparameters for all baselines are favorably configured following the original papers. All methods are implemented with PyTorch 2.0.0 and executed on NVIDIA A100 GPUs.
@@
Evaluation. We evaluate T2I-alignment by GPT-4o and humans. For precise evaluation, we ask GPT-4o and humans with a detailed score rubric (Chen et al., 2024a), which are detailed in Appendix F.
@@ 4.2 Main Results of R2F
Table 4: Performance of R2F combined with different diffusion models (SDXL and SD3.0) on RareBench.
@@ 4.3 Flexibility across Various Diffusion Models and LLMs
Table 5: Performance of R2F combined with different LLMs (GPT-4 and LLaMA3).
@@ 4.4 Ablation Studies
Table 7: Controllable image generation performance of R2F+ on three benchmarks with multi-object cases. T2I alignment accuracy measured from GPT-4o was reported.
@@ 4.5 Controllable Image Generation Results of R2F+
Table 4.4 shows the superiority of R2F+ for controllable image generation on three benchmarks with multi-object cases. Overall, R2F+ mostly performs the best in terms of T2I alignment accuracy, outperforming even R2F by leveraging more detailed layout-guided image generation process as shown in Figure 10. In addition, Figure 11 visualizes the generated images of R2F+ compared to SD3.0. With the proper layouts and their rare-to-frequent concept guidance generated by LLM, R2F+ consistently shows very controllable image generation results, achieving proper attribute binding on the corresponding region of objects. Meanwhile, R2F+ is not able to accurately synthesize images from prompts with ‘complex’ cases where multi-objects are intertwined in the region. This may be because the training-free layout-guided generation often fails with overlapped bounding boxes (Yang et al., 2024). More analysis with failure cases of R2F+ can be found in Appendix J.
@@ Appendix C LLM Instruction and Algorithm Pseudocode for R2F+
To obtain the object mask $M^{i}$ for object $i$ from its object-wise image, we follow a two-step process. First, an object detection model such as Grounding DINO (Liu et al., 2023) identifies the bounding box corresponding to $\mathbf{c}^{i}_{R}$. This bounding box may differ from the provided region guide $R^{i}$. Next, a segmentation model such as SAM (Kirillov et al., 2023) generates the segmentation mask based on the bounding box. The resulting mask is resized from the image space to the latent space.
@@ Appendix F Details for Evaluation
Focus on the objects in the image and their attributes (such as color, shape, texture), spatial layout and action relationships.
According to the image and your previous answer, evaluate how well the image aligns with the text prompt: [PROMPT]
Give a score from 0 to 5, according the criteria:
@@ Appendix I Failure Results of T2I-CompBench’s Auto-evaluation Metrics
Figure 15 shows more visualization examples and failure cases of R2F+ on RareBench. As shown in the first row (i.e., Successful Cases), R2F+ produces highly controllable image generation results when the bounding boxes are non-overlapped and assigned with proper sizes. However, as shown in the second row (i.e., Failure Cases), when the bounding boxes are overlapped or too small, which is usually produced by LLM when the prompt is long and complex, it fails to accurately generate visual concepts. For example, the boundaries between objects are blurred, and the detailed attributes disappear, which is often reported by the existing literature of region-guided diffusion (Chen et al., 2024b, Yang et al., 2024).

### v2

@@ Abstract
State-of-the-art text-to-image (T2I) diffusion models often struggle to generate rare compositions of concepts, e.g., objects with unusual attributes. In this paper, we show that the compositional generation power of diffusion models on such rare concepts can be significantly enhanced by the Large Language Model (LLM) guidance. We start with empirical and theoretical analysis, demonstrating that exposing frequent concepts relevant to the target rare concepts during the diffusion sampling process yields more accurate concept composition. Based on this, we propose a training-free approach, R2F, that plans and executes the overall rare-to-frequent concept guidance throughout the diffusion inference by leveraging the abundant semantic knowledge in LLMs. Our framework is flexible across any pre-trained diffusion models and LLMs, and can be seamlessly integrated with the region-guided diffusion approaches. Extensive experiments on three datasets, including our newly proposed benchmark, RareBench, containing various prompts with rare compositions of concepts, R2F significantly surpasses existing models including SD3.0 and FLUX by up to $28.1\%p$ in T2I alignment. Code is available at https://github.com/krafton-ai/Rare-to-Frequent.
@@ 1 Introduction
To thoroughly validate the efficacy of R2F, we present a new benchmark, dubbed RareBench, consisting of diverse and complex rare compositions of concepts. On RareBench and two existing compositionality benchmarks (Rassin et al., 2024, Huang et al., 2023), R2F outperforms state-of-the-art diffusion baselines, including SD3.0, IterComp, and FLUX, by up to $28.1\%p$ in terms of T2I alignment accuracy. Moreover, we show that R2F can generate images of rare concepts that existing models almost always fail to generate even with careful prompt paraphrasing, showcasing the superiority of our framework in unlocking the compositional generation power of diffusion models.
@@ 3.2 Proposed Framework: R2F
Inspired by the theory, we propose a training-free framework, R2F, which leverages LLM to find frequent concepts relevant to rare concepts and use them in diffusion sampling. R2F involves a two-stage process: (i) Rare-to-frequent concept mapping that uses LLM to identify rare concepts and their relevant yet frequent concepts; and (ii) Alternating concept guidance that iteratively uses prompts involving either rare or frequent objects during the sampling process. Unlike Theorem 3.1, in multi-step denoising, we use alternating guidance as the default approach for rare-frequent interpolation (See Section  4.4 for its efficacy over interpolation). For the acceleration method with short sampling steps, we use interpolation (See Appendix K). Figure 4 illustrates the overview of our framework.
@@ 4.1 Experiment Setting
Implementation Details. Our R2F is flexible to an arbitrary combination of diffusion models and LLMs. By default, we use SD3.0 and GPT-4o (Achiam et al., 2023). For every inference, we set the sampling steps $T$ to $50$ and the random seed to $42$. The hyperparameters for all baselines are favorably configured following the original papers. All methods are implemented with PyTorch 2.0.0 and executed on NVIDIA A100 GPUs. See Appendix O for GPU efficiency analysis.
@@
Evaluation. We evaluate T2I-alignment by GPT-4o and humans. For precise evaluation, we ask GPT-4o and humans with a detailed score rubric (Chen et al., 2024b), which are detailed in Appendix F.
@@ 4.2 Main Results of R2F
Table 4: Performance of R2F combined with different diffusion models (SDXL, IterComp, and SD3.0) on RareBench.
@@ 4.3 Flexibility across Various Diffusion Models and LLMs
Table 5: Performance of R2F combined with different LLMs (GPT-4o or LLaMA3).
@@ 4.4 Ablation Studies
Table 7: Controllable image generation performance of R2F+ on three benchmarks with multi-object cases. T2I alignment accuracy measured from GPT4 was reported.
@@ 4.5 Controllable Image Generation Results of R2F+
Table 4.4 shows the superiority of R2F+ for controllable image generation on three benchmarks with multi-object cases. Overall, R2F+ mostly performs the best in terms of T2I alignment accuracy, outperforming even R2F by leveraging more detailed layout-guided image generation process as shown in Figure 10. In addition, Figure 11 visualizes the generated images of R2F+ compared to SD3.0. With the proper layouts and their rare-to-frequent concept guidance generated by LLM, R2F+ consistently shows controllable image generation results, achieving proper attribute binding on the corresponding region of objects. Meanwhile, R2F+ can not accurately synthesize images from prompts with ‘complex’ cases where multi-objects are intertwined in the region. This may be because the training-free layout-guided generation often fails with overlapped bounding boxes (Yang et al., 2024). More analysis with failure cases and image quality scores are in Appendix J and Appendix N.
@@ Appendix C LLM Instruction and Algorithm Pseudocode for R2F+
To obtain the object mask $M^{i}$ for object $i$ from its object-wise image, we follow a two-step process. First, an object detection model such as Grounding DINO (Liu et al., 2023b) identifies the bounding box corresponding to $\mathbf{c}^{i}_{R}$. This bounding box may differ from the provided region guide $R^{i}$. Next, a segmentation model such as SAM (Kirillov et al., 2023) generates the segmentation mask based on the bounding box. The resulting mask is resized from the image space to the latent space.
@@ Appendix F Details for Evaluation
Focus on the objects in the image and their attributes (such as color, shape, texture), spatial layout, and action
relationships. According to the image and your previous answer, evaluate how well the image aligns with the
text prompt: [PROMPT]
Give a score from 0 to 5, according to the criteria:
@@ Appendix I Failure Results of T2I-CompBench’s Auto-evaluation Metrics
Figure 15 shows more visualization examples and failure cases of R2F+ on RareBench. As shown in the first row (i.e., Successful Cases), R2F+ produces highly controllable image generation results when the bounding boxes are non-overlapped and assigned with proper sizes. However, as shown in the second row (i.e., Failure Cases), when the bounding boxes are overlapped or too small, which is usually produced by LLM when the prompt is long and complex, it fails to accurately generate visual concepts. For example, the boundaries between objects are blurred, and the detailed attributes disappear, which is often reported by the existing literature of region-guided diffusion (Chen et al., 2024c, Yang et al., 2024).
@@
Appendix K Accelerated R2F using FLUX with 4-step Inference.
Here, we further propose an accelerated version of R2F integrated with FLUX-schnell which only requires 4 or fewer steps to generate each image. In scenarios with short inference lengths, the efficacy of alternating guidance may diminish. Therefore, while maintaining R2F’s main idea of exposing frequent concepts to the diffusion sampling process, we propose the Composable method (Same as the one we used in Section 4.4) as an alternative.
Configuration. Given a pair of rare-frequent concept prompts, Composable blends the text embeddings of two prompts and uses them as the input for diffusion sampling steps. For the guidance length of 4 steps, we adopted its interpolation configuration as follows: (1) For concepts with a visual detail level from 1 to 3, we applied the composable method only for the first step with the blending factor $\alpha$ of 0.3. (2) For concepts with a visual detail level from 4 to 5, we further applied it until the second step with the decreased blending factor $\alpha$ of 0.3. Only the original rare prompt was exposed in the final third and fourth steps.
Table 15: 4-step inference result of R2F combined with FLUX-schnell.
RareBench Property Shape Texture Action Complex
FLUX 72.5 68.1 49.3 61.2 73.7
[…]
Results. Table 15 shows the 4-step inference result of R2F combined with FLUX-schnell on RareBench single case. Owing to the adopted frequent concept exposure approach based on the Composable method, R2F improves the compositional generation ability of FLUX even with short guidance steps of 4. Figure 16 visualizes the generated images of R2F combined with FLUX. Overall, R2F significantly enhances the T2I alignment of the generated image without compromising the image quality. This indicates the frequent concept exposure idea of R2F can be generalized to the latest acceleration methods with short sampling steps, which leads to a broader impact on many real-world applications where fast inference time is crucial.
Figure 16: Generated images of R2F using FLUX as the backbone with a 4-step denoising process.
Appendix L Visualization Results of R2F when Exposing Frequent Concept until the Last Step.
Figure 17 shows the generated images of R2F when using concept alternating at the beginning steps and using frequent concepts at the last steps. Overall, the generated image tends to align more closely with the frequent prompt rather than the original rare prompt. This is because diffusion is a step-wise denoising process where the generated image depends more on the prompt that has been used lately.
Figure 17: Visualization examples of R2F when exposing frequent concept until the last step.
Appendix M Effect of LAION-400M Dataset for Finding Rare-Frequent Concept Mapping.
Here, we investigate the effect of LAION-400M dataset (Schuhmann et al., 2022), one of the largest open-source image-caption datasets available online, to enhance the performance of R2F in terms of finding more appropriate rare-frequent concept mapping.
Setup. Given a prompt (e.g., “a bearded apple”), if it contains a rare attribute word (“bearded”), we measure the frequency of its next words in the LAION-400M. For example, if there are 100 captions containing “bearded man” in LAION-400M, the frequency of “man” for the attribute “bearded” is calculated as 100. We then integrated these next-word frequencies into the R2F process in two ways: (1) using the most frequent next word as the frequent concept. For each rare attribute word, we extract the most frequent subsequent noun from LAION and directly use it for the noun of the frequent concept. For example, if “man” appears most frequently after “bearded”, we use “bearded man” as the frequent concept for the original rare concept such as “bearded apple”. (2) providing Top20 next word frequency in LLM prompt. In this case, we extract the top 20 frequent next words from LAION-400M, and provide them to the LLM prompting for identifying rare-to-frequent concept mapping with the following instruction; “…When finding frequent concepts for extracted rare concepts, please consider the words that appeared most frequently after the attribute word of the rare concept in the LAION image caption dataset. The list of the top 20 words is as follows and is in the format of (‘next word’, ‘count’). EXAMPLES…”.
Table 16: Effect of LAION-400M for finding rare-frequent concept mapping in R2F.
Models SD3 R2F R2F w/ (1) R2F w/ (2)
$\text{{RareBench}{}}_{\text{property}}$ 49.4 89.4 81.3 85.9
Result. Table 16 shows the effect of LAION-400M for finding rare-frequent concept mapping. While we expose the word frequency information from LAION-400M into R2F, the performance of these variants is not higher than the original R2F. This may be because the LAION captions are low-quality (e.g., captions are mostly alt texts crawled from the web) and recent models such as SD3.0 are trained on more high-quality undisclosed image-caption datasets (Betker et al., 2023), which potentially diverges from the distribution of LAION captions. For example, the top 5 subsequent words following "bearded" in LAION-400M are (‘man’, 8772), (‘dragon’, 5996), (‘collie’, 3573), (‘iris’, 2153), and (‘dragons’, 1087), showing a discrepancy with common sense knowledge. Nevertheless, all the variants outperform SD3.0, which indicates the effectiveness of the fundamental idea of R2F’s frequent concept exposure.
Appendix N Quantitative Image Quality Analysis of R2F+.
Metrics. For quantitative image quality analysis of R2F+, we use three popular image quality scores, including LAION-aesthetic (Schuhmann et al., 2022), PickScore (Kirstain et al., 2023), and ImageReward (Xu et al., 2024).
Table 17: Aesthetic scores of generated images by R2F+ on RareBench multi-object case.
Scores LAION-aesthetic PickScore ImageReward
R2F 3.980$\pm$0.361 0.226$\pm$0.009 0.626$\pm$0.029
R2F+ 3.887$\pm$0.353 0.222$\pm$0.009 0.609$\pm$0.033
Result. Table 17 compares the image quality scores of R2F+ with R2F. Overall, there is no significant difference between the two models in terms of image quality scores, indicating that R2F+ can achieve better spatial composition without much comprising the image quality compared to R2F.
Appendix O GPU Time and Memory Analysis.
Table 18: GPU time and memory required to generate an image of prompt “a horned lion and a wigged elephant”.
Models SD3 R2F R2F+
Peak Memory (GB) 31.52 31.76 35.08
GPU Time (sec) 20.04 20.60 72.37
To further show the practicality of our algorithms, we provide the GPU time and memory analysis on a single NVIDIA 40GB A100 GPU. We measure the GPU time and memory required to generate a prompt “a horned lion and a wigged elephant”, which consists of two rare concepts “horned lion” and “wigged elephant”. The results are shown in Table 18 (only time and memory taken for diffusion sampling steps are presented).
Peak Memory. While R2F+ involves several latent and gradient computations, there is no significant difference in peak memory compared to R2F since it follows a sequential process. R2F requires approximately 31GB of peak memory, and R2F+ requires approximately 35GB of peak memory where an additional 4GB mostly comes from the gradient computations in cross-attention control.
GPU Time. The time taken for R2F+ is approximately 72 sec, which can be decomposed as 1) masked latent generation via object-wise R2F, which takes around 42 sec, and 2) region-controlled concept guidance, which takes around 30 sec. Specifically, for the process of 1), the generation of each object takes around 20 sec (same as R2F) with an additional 1 sec for masking, resulting in a total of 42 sec for two objects. For the process of 2), the majority of the increased computation time compared to R2F is attributed to attention control, which adds around 10 sec. Consequently, for a prompt with $N$ objects, the time complexity of R2F+ is expected as $N*(T+1)+(T+10)$, where $T$ is the inference time of R2F or SD3.
Appendix P Broad Applications.
Rare concept composition is essential in various applications that require the creation of creative content, such as designing characters and posters for comics, movies, and games. Creators in these domains should often produce content that has never existed, such as characters with elemental faces (e.g., fire or water), pirate ship-shaped spaceships, or trumpet-like guns. Therefore, rare concept composition could be considered a mainstream area for these creators.
Furthermore, the idea of frequent concept guidance can potentially be extended to other modalities, such as text-to-speech (TTS) (Liu et al., 2023a, Lyth & King, 2024, Lee et al., 2024) and text-to-music (Copet et al., 2024). For instance, TTS models have concept categories including speaker, intonation, and emotion. When generating speech such as “an angry Trump speaking Catalan”, we might expose frequent concepts such as “an angry Spanish speaking in Catalan” to improve the composition performance of diffusion-based TTS models.
R2F+ is useful in many applications where the layout-aware composition is essential. (1) Layout-aware image/poster design. When a user creator wants to place an object in a specific position within an image for poster design, R2F is insufficient because it cannot adjust absolute positions. In such cases, R2F+ becomes essential. (2) Data synthesis for enhancing spatial understanding of foundation models. Recent multi-modal LLMs (e.g., LLaVA (Liu et al., 2024, Park et al., 2024)) and pre-trained VLMs (e.g., CLIP (Radford et al., 2021)) are known to exhibit weaknesses in spatial understanding (Chen et al., 2024a), so several studies have attempted to enhance their performance with spatiality-aware image synthesis (e.g., generating images that accurately captures spatial information in text prompts). R2F+ has the potential to enhance the performance of these foundation models by serving as a data synthesis method, when spatial composition is more critical than image quality.